OpenStudy (highschoolmom2010):

The polynomial x^3 + 5x^2 -57x -189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions?

ganeshie8 (ganeshie8):

I think you need to factor the given polynomial

ganeshie8 (ganeshie8):

each factor represents each of the 3 dimensions

OpenStudy (princeharryyy):

x^2(x+3) + 2x(x+3)-63(x+3) = 0 (x+9)(x-7)(x+3)=0 @highschoolmom2010

OpenStudy (princeharryyy):

the three dimensions are in the equation . hope it helps @highschoolmom2010

OpenStudy (princeharryyy):

15,21,8 @highschoolmom2010

OpenStudy (highschoolmom2010):

I don't understand..

OpenStudy (princeharryyy):

x+3 = 15 => x=12 son one side x+9 = 21 second x-7 =12-7 =5 and third x+3 = 15 already given @highschoolmom2010

OpenStudy (highschoolmom2010):

I just don't understand what is being done.

OpenStudy (princeharryyy):

see the box is a cubic structure as given in the ques. now the cubic box volume is given by l*b*h where l, b, h are the length, breadth and height of the cubic structure now if u can write the polynomial in such a way that is in the form of l*b*h here which comes out to be (x+9)*(x-7)*(x+3) in which one is length (x+9) one is breadth (x-7) and the other is height (x+3) u will get the volume of the cubic structure . I hope u'll get it this time @highschoolmom2010